OpenStudy (anonymous):

PLEASE HELP!! iM SO LOST! ITS Algebra 1 standard, 9th Grade (PLEASE SHOW WORK, THANK YOU) -(w+5)= -14

OpenStudy (anonymous):

-(w+5) = -14 -w -5 = -14 -w = -19 w = 19

OpenStudy (anonymous):

... -w = -9 w = 9 sorry

OpenStudy (anonymous):

No its fine! THANK YOU THOUGH!! (:

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

some mistake here i think

OpenStudy (anonymous):

\[-(w+5)= -14 \] \[-w-5=-14\] \[-w=-14+5\] \[-w=-9\] \[w=9\]